OpenStudy (anonymous):

Brad sold candy bars and cookies for a fundraiser at school. Candy bars sold for $3 and cookies sold for $5. He sold a total of 20 items and made $76. How many candy bars did Brad sell?

OpenStudy (anonymous):

Candy bars = B = 3 Cookies = C = 5 Max 20 Items x+y=20 XB + YC = 76 put the values for B and C in 3X + 5Y = 76 X + Y = 20 if X + Y = 20 then X = 20-y insert into 3X + 5Y = 76 3(20-Y) + 5Y = 76 expand brackets 60-3Y + 5Y = 76 solve the rest of this equation and then put your answer into X + Y = 20 and you will have your answers